# Assessing Vulnerability in Smart Contracts: The Role of Code Complexity Metrics in Security Analysis

### v1

@@ VII Experimental Results
In this results section, we give answers to each research question asked.
@@
Among all the coefficients calculated for each group, the p-values for the neutral group were consistently below 0.05, indicating a high level of confidence in the results. However, exceptions were noted in the relationships between CLOC and NOD, DIT and NOD, and NOA and NOD, with p-values of 0.50, 0.09, and 0.09, respectively. Unfortunately, for the vulnerable group, the p-values were higher than 0.05 in many instances. Given the high number of cases, we only report those that are important for our use, specifically when a high correlation is observed. For a full report, you can look at our implementation’s source code which is publicly available at https://zenodo.org/records/13364187
@@
RQ2: Is there a correlation between each complexity metric and the existence of vulnerability? We aimed to determine whether certain complexity metrics are associated with the presence of vulnerabilities in individual contracts. This part also utilizes the Spearman Correlation Coefficient but without separately grouping the codes into vulnerable and neutral. In our study, all of the calculated p-values for the correlations between complexity metrics and vulnerability indicators were less than 0.05. The result of the Spearman correlation coefficient for each metric is represented in Table V. To analyze the result more comprehensively, figure 4 illustrates the result in a bar plot.
@@
TABLE V: Correlations between complexity metrics and existence of Vulnerabilities
@@
Table VI illustrates the comparison of correlation outcomes between complexity metrics in smart contracts and the same metrics in traditional programming languages. Only the metrics consistent with previous research are included. The table findings indicate that while in other languages these metrics exhibited a monotonic relation with vulnerability presence, this relationship was not consistently observed in the context of smart contracts.
TABLE VI: Comparing the correlation values of complexity and existence of vulnerability between smart contracts (SC) and traditional programs (TP)
@@
RQ3: Are the metrics different in vulnerable and neutral codes? We sought to investigate whether the complexity metrics can effectively distinguish between functions with reported vulnerabilities and those without reported vulnerabilities. To assess whether the metrics in each group (vulnerable and neutral) can effectively discriminate between vulnerable and neutral codes, we employed the paired t-test. This statistical test was conducted by comparing the metric values of the vulnerable group with their corresponding values in the neutral group. A p-value greater than 0.05 indicates that the null hypothesis is true, suggesting that the metric’s values are derived from the same distribution and cannot effectively discriminate between vulnerable and neutral codes. On the other hand, a p-value less than or equal to 0.05 suggests that there are significant differences in the metric values between the two groups, indicating the potential of the metric to discriminate vulnerable codes from neutral codes. The p-values and the results are available in Table VII
TABLE VII: The paired t-test between each metric’s values in vulnerable and its corresponding values in neutral codes
@@ VIII Discussion
The dataset used is newly created but only comprises 16,239 contracts, which might be considered small for our 21 metrics. Although there are several large smart contract datasets containing new source codes [45], a sufficiently large and new dataset labeled with vulnerabilities has yet to be developed. Given these limitations, caution must be exercised when generalizing the results of this dataset to the current state of Solidity and its vulnerabilities. Future research endeavors should strive to include larger datasets, up-to-date analysis tools, and labeling approaches that align with the latest versions of Solidity. By doing so, we can ensure that our conclusions and evaluations accurately reflect the most relevant and current vulnerabilities present in Smart Contracts.
@@
IX Conclusion
@@
Although the Spearman rank correlation coefficients indicated low individual correlations between each complexity metric and vulnerability, our machine learning models still achieved good results. This suggests that no single metric can effectively distinguish between vulnerable and invulnerable contracts on its own. However, when combined, the metrics collectively provide practical information and enhance the performance of the models. This highlights the importance of considering multiple complexity metrics together rather than relying on a single metric.
This study provides practical evidence of the effectiveness and relevance of complexity metrics in predicting vulnerabilities in Solidity smart contracts and analyses the power of each metric to help researchers understand their individual influence. These metrics can enhance existing models and methodologies by providing additional insights into code vulnerability assessment.
@@
The implementations, source code, data, and experimental results are publicly available at https://zenodo.org/records/13364187

### v2

@@ VII Experimental Results
In this results section, we give answers to each research question asked. For a full report, you can look at our implementation’s source code which is publicly available at Zenodo [^4] https://zenodo.org/records/13364187 and GitHub [^5] https://github.com/MasoudJTehrani/SC-Complexity-Vulnerability
@@
Among all the coefficients calculated for each group, the p-values for the neutral group were consistently below 0.05, indicating a high level of confidence in the results. However, exceptions were noted in the relationships between CLOC and NOD, DIT and NOD, and NOA and NOD, with p-values of 0.50, 0.09, and 0.09, respectively. Unfortunately, for the vulnerable group, the p-values were higher than 0.05 in many instances. Given the high number of cases, we only report those that are important for our use, specifically when a high correlation is observed.
@@
RQ2: Is there a correlation between each complexity metric and the existence of vulnerability? We aimed to determine whether certain complexity metrics are associated with the presence of vulnerabilities in individual contracts. This part also utilizes the Spearman Correlation Coefficient but without separately grouping the codes into vulnerable and neutral. In our study, all of the calculated p-values for the correlations between complexity metrics and vulnerability indicators were less than 0.05. The result of the Spearman correlation coefficient for each metric is represented in Table IV. To analyze the result more comprehensively, figure 4 illustrates the result in a bar plot.
@@
TABLE IV: Correlations between complexity metrics and existence of Vulnerabilities
@@
Table V illustrates the comparison of correlation outcomes between complexity metrics in smart contracts and the same metrics in traditional programming languages. Only the metrics consistent with previous research are included. The table findings indicate that while in other languages these metrics exhibited a monotonic relation with vulnerability presence, this relationship was not consistently observed in the context of smart contracts.
TABLE V: Comparing the correlation values of complexity and existence of vulnerability between smart contracts (SC) and traditional programs (TP)
@@
RQ3: Are the metrics different in vulnerable and neutral codes? We sought to investigate whether the complexity metrics can effectively distinguish between functions with reported vulnerabilities and those without reported vulnerabilities. To assess whether the metrics in each group (vulnerable and neutral) can effectively discriminate between vulnerable and neutral codes, we employed the paired t-test. This statistical test was conducted by comparing the metric values of the vulnerable group with their corresponding values in the neutral group. A p-value greater than 0.05 indicates that the null hypothesis is true, suggesting that the metric’s values are derived from the same distribution and cannot effectively discriminate between vulnerable and neutral codes. On the other hand, a p-value less than or equal to 0.05 suggests that there are significant differences in the metric values between the two groups, indicating the potential of the metric to discriminate vulnerable codes from neutral codes. The p-values and the results are available in Table VI
TABLE VI: The paired t-test between each metric’s values in vulnerable and its corresponding values in neutral codes
@@ VIII Discussion
Based on the findings of this study, we can present how complexity metrics can be integrated into existing vulnerability detection tools. Also, we discuss how this study’s findings might influence smart contract best practices and developer guidelines. Lastly, we talk about the future directions, and how to extend the study of complexity in smart contracts.
VIII-A Integrating Complexity Metrics
Complexity metrics can be integrated into existing vulnerability detection tools by leveraging their ability to distinguish between vulnerable and neutral smart contract codes. This study’s findings suggest several practical approaches for enhancing vulnerability detection tools:
  • Feature Optimization: This study identifies redundancy among complexity metrics, such as strong correlations between SLOC, WMC, and NF. By selecting only the most relevant metrics, vulnerability detection models can be made more efficient, reducing computational overhead without sacrificing accuracy.
  • Risk Assessment Models: The correlation between certain complexity metrics and vulnerability presence, even if weak, suggests that these metrics can serve as supplementary indicators in risk assessment models. Metrics such as CBO (Coupling Between Objects) and NOI (Number of Invocations) show relatively higher correlations with vulnerabilities and can be weighted accordingly in predictive models.
  • Automated Analyzers: Smart contract security tools can incorporate discriminative complexity metrics identified in the study (e.g., SLOC, WMC, NOI) as warning signals, flagging contracts with high complexity as more likely to contain vulnerabilities. This aligns with best practices for keeping smart contracts simple and auditable.
  • Static Analysis Enhancements: Existing static analysis tools can integrate complexity-based heuristics to prioritize contract segments that require deeper analysis. Since vulnerability likelihood increases with complexity, tools can focus audits on high-complexity regions to enhance security assessments.
VIII-B Best Practices
Regarding best practices and developer guidelines, the study’s findings highlight the following key takeaways:
  • Simplicity is Key: Higher complexity is generally associated with a greater likelihood of vulnerabilities. Developers should aim to minimize unnecessary functions, control structures, and deep nesting to reduce the risk of security flaws.
  • Code Documentation Matters: Unlike traditional software, smart contracts with higher comment density (CLOC) were found to be less vulnerable, suggesting that well-documented contracts may result in better observation and have fewer security issues.
  • Selective Use of Inheritance: Metrics related to inheritance (e.g., NOA, DIT) were found to be discriminative, meaning that complex inheritance structures might introduce subtle vulnerabilities. Developers should keep inheritance hierarchies shallow and transparent.
  • Tool Support for Complexity Awareness: Smart contract development tools and IDEs should integrate complexity visualization and alerts to help developers write safer code by identifying potentially risky patterns early in development.
By utilizing these takeaways in vulnerability detection tools and developer guidelines, the security of smart contracts can be significantly improved, leading to more reliable and robust decentralized applications.
VIII-C Future Work
@@
IX Threat To Validity
IX-A External Threats
Given the critical role of datasets in this type of research, any changes in the dataset or labeling could lead to minor variations in experimental outcomes. To mitigate this, this paper utilizes a widely recognized and frequently used dataset in smart contract research. As a result, while some variation may occur with a different dataset, the overall analysis remains valuable and relevant.
The smart contract codes in the dataset were collected using a standard approach from a public resource. Due to the large number of contracts, bias toward a specific Solidity version was minimized. However, since the syntax of older Solidity codes is largely different, collecting older smart contracts and replicating this study on older versions of Solidity may lead to different results.
IX-B Internal Threats
This paper employed statistical hypothesis tests, including paired t-tests, to identify significant differences between paired data sets. Since these tests rely on several assumptions, their reliability—particularly for RQ3—may be affected if those assumptions are not met in a different setup. To mitigate this, these methods were selected based on prior research and the characteristics of the dataset. This also ensured the consistency with studies used for comparison.
Similar to the selection of statistical tests, the thresholds and intervals were chosen based on best practices and the dataset. It is reasonable to expect that using different values for these variables could lead to varying results.
X Conclusion
@@
Although the Spearman rank correlation coefficients indicated low individual correlations between each complexity metric and vulnerability, our machine learning models still achieved good results. This suggests that no single metric can effectively distinguish between vulnerable and invulnerable contracts on its own.
This study provides practical evidence of the effectiveness and relevance of complexity metrics in predicting vulnerabilities in Solidity smart contracts and analyses the power of each metric to help researchers understand their individual influence. Furthermore, this study discusses the methods of how complexity metrics can be integrated into existing vulnerability tools as well as how developers can influence these findings and implement a safer code.
@@
The implementations, source code, data, and experimental results are publicly available at https://zenodo.org/records/13364187 and https://github.com/MasoudJTehrani/SC-Complexity-Vulnerability